How do I solve with x as power?
2^x+4 = 4(16^x)

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

How do I solve with x as power?
2^(x+4) = 4(16^x)
---
2^(x+4) = 4(2^4x)
2^(x+4) = 2^2(2^4x)
2^(x+4) = 2^(4x+2)
x+4 = 4x+2
3x = 2
x = 2/3
==============
Cheers,
Stan H.

Answer by solver91311(24713)



Use logarithms. Any base will do in the final analysis, but the arithmetic will be neater in this case if you use base 2.

First, I will make the assumption that you mean:



as opposed to:



First note that

Hence your equation becomes:



Now take the base 2 logarithm of both sides:



Use the following laws of logarithms to simplify the equation:







Next note that

So:



But also recall that:



So we can say:



and then:



Which you can solve by ordinary means.

John

My calculator said it, I believe it, that settles it